The present value (at n=0) of the following cash flow stream is $5,979.04 when discounted at 12% annually. That stream consists of payments made at the end of each of four periods. One of the payments, the second one, is missing. The others are $1,000 at n=1, $2,000 at n=3, and $2,000 at n=4. What is the value of the missing cash flow? Please show using excel.
Answer #1

Present Value of Uneven cash flow can be calculated as

CF0 + CF1/(1+r)^1 + CF2/(1+r)^2 + ...... + CFn/(1+r)^n

5979.04 = 1000/1.12^1 + CF2/1.12^2 + 2000/1.12^3 + 2000/1.12^4

= 1000/1.12 + CF2/1.2544 + 2000/1.404928 + 2000/1.57351936

= 892.857142857 + CF2/1.2544 + 1423.56049563 + 1271.03615681

=CF2/1.2544 + 3587.4537953

CF2/1.2544 =5979.04 -3587.4537953

= 2391.5862047

CF2 = 2391.5862047*1.2544

= 3000